OpenStudy (roseliberty):

Please help Why there is no solution for log3(-4) ?

OpenStudy (mathstudent55):

The domain of the log function is non-negative numbers. The log of zero and the log of negative numbers are undefined.

OpenStudy (mathmale):

Review the concept of domain. Mathstudent55 (above) has it right.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \log_{3}(-4)=x \hspace{.33em}\\~\\ & 3^{x}=-4 \hspace{.33em}\\~\\ & \text{which is not possible} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathstudent55):

Think of what log means. \(\log_b x = y\) means \(b^y = x\) The log of a number is the exponent you need to raise the base to, to get the number. The \(\log_{10} 100 = 2\) since \(10^2 = 100\) The log of a negative number could be written as this: \(\log_{10} -100 = y\) means that \(10^y = -100\) There is no exponent, y, that you can raise 10 to, to get -100.